OpenStudy (anonymous):

If H=24 units and R=8 units , what is the volume of the cone shown below ? use 3.14 for pie A.1,607.68 cubic units B.602.88 cubic units C.4,823.04 cubic units D.200.96 cubic units

OpenStudy (anonymous):

http://adf.ly/7C53

OpenStudy (anonymous):

that makes me download something @Icedragon50

OpenStudy (anonymous):

oh ok then don't click on it

OpenStudy (anonymous):

i didnt do u know the answer @Icedragon50

OpenStudy (anonymous):

@Nnesha ?

Nnesha (nnesha):

\[\huge\rm v =\frac{ 1 }{ 3 }\pi r^2 h\] radius = r height = h so replace variables by their values and solve

OpenStudy (anonymous):

@amorfide ? help plz medal ?

Nnesha (nnesha):

i already gave you the formula please replace variables and use calculator

OpenStudy (amorfide):

didn't we do the same type of question last time lol

OpenStudy (anonymous):

yes lol @amorfide its homework

Nnesha (nnesha):

okay try to do it that will help you on a test

OpenStudy (anonymous):

i did i need help @Nnesha

OpenStudy (amorfide):

okay you know that h=24 r=8 and pi is delicious so substitute your values into \[V=\pi(r^{2})\frac{ h }{ 3 }\] to get \[V=\pi(8^{2})\frac{ 24 }{ 3 }\]

OpenStudy (amorfide):

you can do this I am sure

OpenStudy (anonymous):

i got b @amorfide

OpenStudy (amorfide):

now you are just guessing

OpenStudy (amorfide):

put it into your calculator

OpenStudy (amorfide):

\[8^{2}=64\] so you have \[V=\pi(\frac{ 64 \times 24 }{ 3 })\] but you know \[\frac{ 24 }{ 3 }=8\] so you have \[V=\pi(8^3)\]

OpenStudy (anonymous):

do u know the answer plz running out of time @amorfide

OpenStudy (amorfide):

use your calculator

Nnesha (nnesha):

plz use calculator

OpenStudy (amorfide):

use pi=3.14

Nnesha (nnesha):

running out of time ?? :O i guess not for h.w

OpenStudy (amorfide):

lmao

OpenStudy (anonymous):

it is so stop d/e plz @Nnesha

Nnesha (nnesha):

okay jw so use calculator and let us know what you get

OpenStudy (anonymous):

i ran out of time now its to late ! @Nnesha

OpenStudy (amorfide):

you should've used a calculator...

Nnesha (nnesha):

^

OpenStudy (anonymous):

its okay @amorfide